OpenStudy (anonymous):

In the circle below, AB and CD are chords intersecting at E. If AE= 5, BE=12, and CE=6, what is the length of DE?

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

you may use this - triangle AED is similar to triangle CEB thus, AE * EB = CE * DE 5 * 12 = 6 * DE DE = ?

OpenStudy (anonymous):

got it thank you {:

ganeshie8 (ganeshie8):

great :) yw!